OpenStudy (anonymous):

What is the value of f[g(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? (Type only the value of your final answer, no variables.)

myininaya (myininaya):

Do you know how to find g(5)?

OpenStudy (anonymous):

hello xD

OpenStudy (anonymous):

g(5)=5 isnt it?

OpenStudy (anonymous):

2*5-5=5

myininaya (myininaya):

yep ! now since g(5)=5 then we need to find f(g(5))=f(5) we need f(5) and we are done! :)

OpenStudy (anonymous):

f(x)= 2x+1 2*5+1=11

OpenStudy (anonymous):

and thats where im stuck

myininaya (myininaya):

yep yep :)

myininaya (myininaya):

you are done!

OpenStudy (anonymous):

o i am lol

OpenStudy (anonymous):

so the answer is 11?

myininaya (myininaya):

yep yep :)

OpenStudy (anonymous):

what the heck that was easy lol

OpenStudy (anonymous):

thank you very much

myininaya (myininaya):

Good job you!

myininaya (myininaya):

You deserve a medal! :)

OpenStudy (anonymous):

o.0 what do i do with these

OpenStudy (anonymous):

ty btw

OpenStudy (anonymous):

And also i have another question if you dont mind helping

myininaya (myininaya):

Well if you like OpenStudy enough maybe you can stay after you are done getting a through your math courses and make your smartscore awesome. This MAY help you with future employers or getting into some schools.

OpenStudy (anonymous):

o cool

OpenStudy (anonymous):

should i post a new question for my other one or just ask it here

myininaya (myininaya):

Can you post it as a new question? I may have to leave soon. Sorry :(

OpenStudy (anonymous):

ok no problem thank you very much

myininaya (myininaya):

np